Circular No. 6258

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

2EG J0432+2910 S. C. Lundgren, R. S. Foster, and E. Waltman, Naval Research Lab-

oratory; D. J. Thompson, B. L. Dingus, and G. Madejski, Goddard Space Flight Center; J. M. Cordes, Cornell University; and P. L. Nolan, Stanford University, report: “An outburst was detected during Aug. 8–22 from the unidentified EGRET source 2EG J0432+2910, now believed to be an active galactic nucleus (AGN). Preliminary analysis indicates the γ-ray flux above 100 MeV during the flare was enhanced by a factor of 5 over the average flux from this source in 1991–1993 observations (1.4 × 10^{−7} photon cm^{−2} s^{−1}). In the subsequent observation (Aug. 22–Sept. 6), the γ-ray flux returned to near its historic average. The strongest flat spectrum radio source in the 15′ γ-ray error box (11′.4 from the center), 87GB 0430+2859, was identified as a possible counterpart (Dingus et al. 1996, Ap.J., submitted). The 21cm NRAO Very Large Array Sky Survey has mapped this field and finds a source at α = 4^{h}33^{m}37^{s}.80, δ = +29^{o}05′56′′.2 (equinox 2000.0), consistent with an Effelsberg survey position (Neumann et al. 1994, A.Ap. Suppl. 106, 303). Radio monitoring of this source with the Green Bank Interferometer was initiated on Oct. 3. The flux densities of 0.60 ± 0.02 Jy at 2.25 GHz and 0.545 ± 0.005 Jy at 8.3 GHz were 35 percent higher than Effelsberg survey flux-density measurements. In 20 days, the 2.25 GHz flux dropped to 0.47 ± 0.025 Jy with smaller day-to-day variations, while the 8.3-GHz flux density remained constant. An optical counterpart near the limiting magnitude of the Palomar Sky Survey has been identified within 2′′ of the VLA position and within the astrometric uncertainty. The probability of chance superposition within 2′′ in this field is 1 percent. A ROSAT HRI observation made on 1995 Mar. 2-4 reveals an x-ray source within 5′′ of the VLA position, which is consistent with the absolute pointing accuracy. The combination of a γ-ray flare, radio variability in a candidate flat spectrum counterpart, and possible optical and x-ray counterparts provides strong evidence that the source is in the blazar class of radio-loud AGN. This source has the weakest radio flux density of any of the blazars detected by EGRET. Further observations are recommended to determine the redshift, measure optical variability, and monitor the high-frequency (> 10 GHz) radio flux associated with the γ-ray flare.”

COMET 94P/RUSSELL 4 CCD magnitudes: Oct. 21.48 UT, 21.4 (J. V. Scotti, Kitt Peak, 0.91-m

reflector); 26.33, 20.5 (W. Offutt, Cloudcroft, NM, 0.60-m reflector).

1995 November 7 Daniel W. E. Green